OpenStudy (anonymous):

Solve for x: −3(x + 3) = −3(x + 1) − 5.

OpenStudy (amoodarya):

first step distribution of (-3) to (x+3) and(x+1) apply this second step simplifying

OpenStudy (anonymous):

what you mean

OpenStudy (amoodarya):

\[−3(x + 3) = −3(x + 1) − 5\\-3x-3(3) =-3x-3(1)-5\\-3x-9=-3x-3-5\\-3x-9=-3x-8\\\]

OpenStudy (amoodarya):

can you go on ?

OpenStudy (anonymous):

yea dont you do the same to the other side ?

OpenStudy (amoodarya):

yes . if we simplify -3x both sides what happen ?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):

count it..try

OpenStudy (anonymous):

take the x to left hand side

OpenStudy (anonymous):

that cancels out right ?

OpenStudy (anonymous):

why cancel?

OpenStudy (amoodarya):

-3x+3x-9=-8 0x-9=-8 -9=-8 now what's the answer ?

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

means the x is undefined. u prove that x cannot be solve.